# A Survey on 3D Gaussian Splatting Applications: Segmentation, Editing, and Generation

### v3

@@ 3.2.4 Other Editing Tasks
$\bullet$ Drag. MVDrag3D [186] employs a multi-view diffusion model as a robust generative prior to achieve consistent drag editing across multiple rendered views. DYG [116] uses 3D masks and control point pairs to define the target editing area and drag direction, leveraging the advantages of implicit tri-plane representation to establish a geometric framework for the editing results.
@@ 3.3.1 Object-level Generation
$\bullet$ Latent-space Optimization. These methods learn a latent space from conditional inputs such as RGB images, depth maps, or point clouds via an encoder, which is then decoded into 3D Gaussians. (a) Text-to-3D. Wizadwongsa et al. [223] employ a feed-forward reconstruction encoder with pre-trained models to reduce training cost, refining unstructured latents via post-processing. They introduce a 2D perceptual rendering loss and a multi-stream transformer correction flow for efficient, high-quality text-conditioned generation. GaussianAnything [224] proposes a scalable framework with a point cloud-structured latent space, supporting multi-modal inputs, geometry–texture disentanglement, and 3D-aware editing. Atlas-Gaussians [225] use a patch-based encoder–decoder with UV sampling and a transformer decoder for efficient, high-fidelity Gaussian shape generation. Turbo3D [226] integrates a four-step, four-view latent diffusion generator with a feed-forward Gaussian reconstructor, employing a dual-teacher scheme for view consistency and photorealism while avoiding pixel-space decoding. (b) Image-to-3D. AGG [227] leverages a pre-trained DINOv2 [46] encoder and dual transformers to decode Gaussian position and texture fields for joint optimization. HumanSplat [228] encodes multi-view inputs via a VAE and applies a latent reconstruction transformer interacting with structured human models. Zou et al. [229] design a hybrid triplane Gaussian latent representation from images and point clouds, decoded into 3D objects with transformer-based point and triplane decoders.
@@ 3.3.2 Scene-level Generation
$\bullet$ Optimization Diffusion Models. (a) Text-to-3D. Prometheus[269] sequentially trains GS-VAE and MV-LDM for direct text-to-3D scene generation. VideoRFSplat [270] couples a pose generation model with a pretrained video generator via communication blocks to produce multi-view images and camera poses. (b) Image-to-3D. Wonderland [271] integrates diverse camera trajectories via a dual-branch conditioning mechanism for view-consistent latents. GGS [272] combines 3D representations with latent video diffusion and a custom decoder to synthesize scenes from feature fields. VideoScene [273] distills video diffusion through a 3D-aware leap flow strategy for one-step generation. SceneSplatter [274] adopts a momentum-based paradigm to balance generative priors and existing scene information during generation.
4 Performance Comparison
@@
4.1 Performance Benchmarking: 3DGS Segmentation
$\bullet$ Datasets. Appendix Table A.1 summarizes the key features of the datasets for 3DGS segmentation.
@@ 4.1 Performance Benchmarking: 3DGS Segmentation
TABLE I: Quantitative 3D instance segmentation on Replica[275] and LERF-Mask[22]. Following [20], we report mIoU and F-score on Replica, and mIoU and mBIoU on LERF-Mask.
@@
TABLE II: Quantitative 3D interactive segmentation on NVOS [276] and SPIn-NeRF[67] in terms of mIoU and mAcc.
@@
TABLE III: Performance comparison of open-vocabulary 3D semantic segmentation on ScanNet [244] benchmark. The evaluation metrics are mIoU and mAcc, respectively.
@@
$\bullet$ Metrics. Common evaluation metrics in the segmentation domain are outlined below:
Mean Intersection over Union (mIoU) measures segmentation accuracy by computing the average IoU across all classes, where IoU is the ratio between the intersection and the union of predicted and ground truth regions for each class.
Mean Accuracy (mAcc) measures the proportion of correctly predicted pixels or instances over the total, reflecting the overall correctness of the segmentation or classification results.
Mean Boundary Intersection over Union (mBIoU) computes the average IoU over narrow boundary regions, emphasizing boundary precision compared to standard mIoU.
Panoptic Quality (PQ) calculates panoptic segmentation performance by recognition quality (RQ) and segmentation quality (SQ).
F-score measures the balance between precision and recall, computed using an IoU threshold of 0.5.
$\bullet$ Results. Following the taxonomy in Sec.2.1.1, we report results across four representative segmentation settings. For 3D instance segmentation (Table I), Unified-Lift [20] achieves the best overall performance on both the Replica [275] and LERF-Mask [22] datasets, demonstrating the effectiveness of consistency-aware end-to-end learning. In the setting of 3D interactive segmentation (Table II), GaussianCut [29] performs best on NVOS [243], while Click-Gaussian [24] attains the highest mIoU on SPIn-NeRF [67]. Notably, GaussianCut is a training-free method that flexibly supports both 2D clicks and text prompts as input. For open-vocabulary 3D semantic segmentation (Table III), PanoGS [118] and InstanceGaussian [31] achieve the best performance under point-based evaluation protocols, whereas CAGS [108] performs best under 3D Gaussian-based evaluation. In the open-vocabulary 2D semantic segmentation setting, VLGaussian [28] reaches a state-of-the-art mIoU of 97.1% on the 3D-OVS benchmark [63], significantly outperforming existing approaches in Table IV. Table V further presents results for two subtasks: open-vocabulary 2D localization (detection) and semantic segmentation. FMLGS [102] achieves the best performance on the segmentation task, while VLGaussian [28] leads in localization. Across these settings, most methods leverage powerful vision-language foundation models such as CLIP and SAM to enhance open-vocabulary capabilities.
TABLE IV: Quantitative open-vocabulary 2D semantic segmentation on 3D-OVS [63] dataset in terms of mIoU.
@@
TABLE V: Quantitative results on LERF-OVS [26] Dataset. We report the open-vocabulary localization accuracy (mAcc) and 2D semantic segmentation (mIoU).
@@
TABLE VI: Performance comparison of 3DGS editing on Mip-NeRF360 [245] and IN2N [75] datasets. The similarity for CLIP Text-Image Direction,CLIP Text-Image, CLIP Image-Image, DINO are denoted as $\text{CLIP}_{dir}$, $\text{CLIP}_{T2I}$, $\text{CLIP}_{I2I}$, and $\text{DINO}_{sim}$, respectively.
@@
TABLE VII: Performance comparison of 3DGS style transfer on Mip-NeRF360 [245] dataset, evaluated by LPIPS and RMSE.
@@
4.2 Performance Benchmarking: 3DGS Editing
$\bullet$ Datasets. We summarize the key features of the datasets in Appendix Table A.2, with detailed descriptions provided below.
@@ 4.2 Performance Benchmarking: 3DGS Editing
$\bullet$ Metrics. Commonly used evaluation metrics in the editing domain are listed below:
CLIP Text-Image Direction Similarity. CLIP [47] extracts two paired sets of text-image embeddings (original and modified), and computes their directional shifts via cosine similarity.
CLIP Text-Image Similarity. CLIP [47] measures text-image similarity by encoding text and images into embeddings separately and then computing their cosine similarity.
CLIP Image-Image Similarity. CLIP [47] computes image-image similarity by encoding edited image and subject image into embeddings separately and then computing their cosine similarity.
Frechet Inception Distance (FID) calculates the similarity between ground truth and generated images by comparing their Inception-v3 [291] feature distributions using the Frechet distance.
DINO Similarity. DINO [292] extracts features from the edited scene and reference image, then computes their average visual similarity to evaluate image-image alignment.
Peak Signal-to-Noise Ratio (PSNR) measures image quality by comparing the maximum pixel intensity to the MSE between images, converted to decibels.
Structural Similarity (SSIM) measures image similarity by comparing brightness, contrast, and structure.
RMSE. It is computed as the root mean squared pixel-wise difference between the edited image and ground truth.
Learned Perceptual Image Patch Similarity (LPIPS) measures perceptual similarity by comparing deep features of two images.
Aesthetic score. Aesthetic Predictor V2.5 [293] is a SigLIP-based model that rates image aesthetics from 1 to 10.
User study. It evaluates results by collecting and analyzing feedback from human participants across customized dimensions such as preference in comparative results, text-image alignment, and view consistency.
@@
$\bullet$ Results. For 3D image editing tasks, the most commonly used datasets are Mip-NeRF360 [245] and IN2N [75]. We adopt these two datasets as benchmarks to record the performance of various methods. However, as most existing 3D editing methods are evaluated on different scene subsets, it remains difficult to perform a comprehensive and fair comparison under a unified protocol. Given that IN2N [75] is widely adopted as a baseline, we report the relative gains of each method over IN2N to enable an indirect yet relatively fair comparison, as summarized in Table VI. Overall, GaussianVTON[170] demonstrates the best performance across multiple metrics. For the style transfer task, we use Mip-NeRF 360 [245] as the benchmark and compare the results of different methods in Table VII, where SemanticSplatSty.[182] achieves the best overall performance.
@@
TABLE VIII: Quantitative 3DGS generation on GSO [247] dataset. CLIP Image-Image Similarity as $\text{CLIP}_{I2I}$.
@@
TABLE IX: Quantitative 3DGS generation on Objaverse [246] dataset. CLIP Text-Image Similarity as $\text{CLIP}_{T2I}$.
@@
4.3 Performance Benchmarking: 3DGS Generation
$\bullet$ Datasets. We summarize the key features of the datasets in Appendix Table A.3, with detailed descriptions provided below.

### v4

@@ 3.2.4 Other Editing Tasks
$\bullet$ Drag. MVDrag3D [268] employs a multi-view diffusion model as a robust generative prior to achieve consistent drag editing across multiple rendered views. DYG [195] uses 3D masks and control point pairs to define the target editing area and drag direction, leveraging the advantages of implicit tri-plane representation to establish a geometric framework for the editing results. 3DGS-Drag [269] enables intuitive 3D drag editing by combining deformation guidance for consistent geometric changes and diffusion guidance for content refinement.
@@ 3.3.1 Object-level Generation
$\bullet$ Latent-space Optimization. These methods learn a latent space from conditional inputs such as RGB images, depth maps, or point clouds via an encoder, which is then decoded into 3D Gaussians. (a) Text-to-3D. Wizadwongsa et al. [305] employ a feed-forward reconstruction encoder with pre-trained models to reduce training cost, refining unstructured latents via post-processing. They introduce a 2D perceptual rendering loss and a multi-stream transformer correction flow for efficient, high-quality text-conditioned generation. GaussianAnything [306] proposes a scalable framework with a point cloud-structured latent space, supporting multi-modal inputs, geometry–texture disentanglement, and 3D-aware editing. Atlas-Gaussians [307] use a patch-based encoder–decoder with UV sampling and a transformer decoder for efficient, high-fidelity Gaussian shape generation. Turbo3D [308] integrates a four-step, four-view latent diffusion generator with a feed-forward Gaussian reconstructor, employing a dual-teacher scheme for view consistency and photorealism while avoiding pixel-space decoding. (b) Image-to-3D. AGG [309] leverages a pre-trained DINOv2 [123] encoder and dual transformers to decode Gaussian position and texture fields for joint optimization. HumanSplat [310] encodes multi-view inputs via a VAE and applies a latent reconstruction transformer interacting with structured human models. Zou et al. [37] design a hybrid triplane Gaussian latent representation from images and point clouds, decoded into 3D objects with transformer-based point and triplane decoders. SAM 3D [311] generates 3D structures in the latent space and then decodes them back into explicit 3D representations.
@@ 3.3.2 Scene-level Generation
$\bullet$ Optimization Diffusion Models. (a) Text-to-3D. Prometheus[344] sequentially trains GS-VAE and MV-LDM for direct text-to-3D scene generation. VideoRFSplat [345] couples a pose generation model with a pretrained video generator via communication blocks to produce multi-view images and camera poses. (b) Image-to-3D. Wonderland [46] integrates diverse camera trajectories via a dual-branch conditioning mechanism for view-consistent latents. GGS [346] combines 3D representations with latent video diffusion and a custom decoder to synthesize scenes from feature fields. VideoScene [347] distills video diffusion through a 3D-aware leap flow strategy for one-step generation. SceneSplatter [348] adopts a momentum-based paradigm to balance generative priors and existing scene information during generation. FlashWorld [349] introduces dual-mode pretraining on a video diffusion model to enable a multi-view model to operate in both MV-oriented and 3D-oriented modes.
4 Challenges and Outlook
4.1 Open Challenges
$\bullet$ Input Artifacts and Reconstruction Robustness. 3DGS-driven downstream applications rely critically on the quality of the underlying 3D reconstruction. In practice, however, images acquired under varying conditions, camera pose errors, uneven view coverage, and artifacts such as blur and transient objects often lead to inconsistent geometry and appearance. While such imperfections may be tolerable for reconstruction itself, they are amplified in downstream tasks, causing unstable semantics and view-inconsistent editing or generation. Some recent methods mitigate these issues through depth regularization, blur-aware reconstruction, artifact-aware optimization, and more robust photometric modeling [20, 84, 83], but they typically target reconstruction quality rather than downstream robustness. Pipelines that explicitly account for input artifacts and other acquisition uncertainties and propagate reconstruction reliability to downstream modules remain largely unexplored, limiting applicability in real-world settings.
$\bullet$ Multi-View Consistency. A central challenge across segmentation, editing, and generation is maintaining semantic and structural consistency across views. Many methods optimize objectives primarily in image space, enforcing cross-view coherence only implicitly through rendering supervision. This can yield strong single-view accuracy while still exhibiting semantic drift, temporal flicker, or inconsistent editing across viewpoints. Existing methods partially address this issue using cross-view feature aggregation, 3D lifting, geometry-aware attention, or multi-view alignment constraints [100, 109, 104, 180], yet most focus on specific tasks and do not generalize across the full range of downstream applications. Designing scalable, task-agnostic consistency mechanisms that jointly preserve geometric structure and semantic identity across views remains an open problem.
$\bullet$ Data Scarcity and Supervision Quality. Progress in 3DGS downstream applications is strongly constrained by the scarcity of reliable 3D annotations. Since dense 3D labels are expensive to obtain, many pipelines lift pseudo labels from 2D foundation models such as SAM [125] and CLIP [124]. While effective, such supervision inevitably introduces noise, bias, and incomplete coverage, particularly for occluded regions, thin structures, or long-tail categories. Several methods attempt to improve label quality through cross-view voting [105], contrastive consistency learning [111, 171], or hierarchical mask refinement [188], but these remain largely heuristic and scene-specific. Scalable supervision strategies or large-scale 3D data are still lacking.
$\bullet$ Efficiency. Despite their strong performance, many 3DGS downstream pipelines remain computationally expensive because they repeatedly optimize a 3D representation and render it back to 2D for supervision. This tight coupling between 3D representation learning and repeated differentiable rendering makes training and adaptation costly, especially for large scenes, long sequences, and interactive settings. Existing methods have begun to address this bottleneck through feed-forward Gaussian prediction, compact representations, pruning and compression, and pretrained reconstruction priors [27, 36, 37, 47, 48, 49]. However, these methods often trade accuracy or robustness for speed, and their benefits are not yet uniform across segmentation, editing, and generation tasks. The challenge is how to achieve a better balance among efficiency, scalability, and downstream performance.
$\bullet$ Evaluation. Existing evaluation protocols largely inherit 2D or NeRF-era metrics such as PSNR, SSIM, and LPIPS, which focus on image-level fidelity and may overlook properties critical to downstream applications, including cross-view semantic consistency, geometric correctness, edit locality, and controllability. Recent efforts have started to incorporate human studies, MLLM-based assessment, and task-specific measures [350, 351, 352]. However, evaluation remains fragmented across datasets and applications. A major open problem is to develop standardized, 3D-aware evaluation frameworks that jointly assess geometry, semantics, multi-view consistency, and practical usability in a unified manner.
4.2 Future Directions
$\bullet$ Combining 3D Foundation Models. To overcome the robustness and multi-view consistency challenges, emerging 3D foundation models such as VGGT [353] and FLARE [354] offer robust representations by pre-training on diverse 3D tasks (e.g., depth estimation, point cloud reconstruction). These models encode rich geometric priors that can stabilize downstream applications under sparse or artifact-heavy inputs. Incorporating them into 3DGS pipelines to improve generalization is a promising direction.
$\bullet$ Toward Generalist Models. Instead of designing separate models for segmentation, editing, and generation, a promising direction is to develop generalist 3DGS-based architectures that perform multiple scene-level tasks within a unified framework. Such models can facilitate cross-task representation sharing, reduce training redundancy, and improve generalization.
$\bullet$ Synthesizing 3D Data. To mitigate the scarcity of reliable 3D supervision, future research can leverage the generative capabilities of 3DGS to synthesize large-scale, high-quality 3D datasets from abundant 2D image collections. Such synthetic data could significantly enhance training diversity and volume. Combining synthesized data with domain adaptation techniques could further bridge the gap between synthetic and real-world distributions, improving robustness and generalization.
$\bullet$ Large-Scale Feed-Forward Learning. To address the efficiency bottleneck of repeated scene-specific optimization, an increasing number of works are moving toward feed-forward architectures. Future research should explore large-scale training regimes across diverse scenarios and domains, including scaling up data volume, improving data diversity (e.g., egocentric, indoor, outdoor), and designing robust pipelines that generalize in the real world.
$\bullet$ Faithful Evaluation Metrics. Current evaluation protocols often inherit metrics from traditional 2D or NeRF-based tasks, which may not align well with 3DGS-specific goals. Developing reliable and interpretable metrics that go beyond 2D projection accuracy to capture inherent 3D geometric and semantic consistency is essential for tracking progress and guiding future research.
5 Conclusion
This survey systematically explores recent advances in utilizing 3D Gaussian Splatting (3DGS) for downstream application tasks, representing a pioneering effort in summarizing this emerging field. We first introduce essential background knowledge, covering fundamental concepts and foundational models. Subsequently, we comprehensively review over 200 representative methods across several key tasks, including segmentation, editing, generation, and other related application tasks, providing a structured categorization from a technical perspective. In the experimental section, we detail the evaluation metrics and compare methods under a fair and consistent protocol.Lastly, we highlight current challenges and outline promising future directions to inspire continued research efforts toward enhanced high-level downstream tasks.
Appendices
[140][106][355][7][102][356][152][357][144][144][98][358][359][360]
A.1 Performance Comparison
@@
TABLE A.1: Statistics of representative 3DGS segmentation datasets. See Sec. 4.1 for more detailed descriptions.
  Dataset Venue #Scene #Views(Avg.) Characterization
A.1.1 Performance Benchmarking: 3DGS Segmentation
$\bullet$ Datasets. Table A.1 summarizes the key features of the datasets for 3DGS segmentation.
@@ A.1.1 Performance Benchmarking: 3DGS Segmentation
TABLE A.2: Quantitative 3D instance segmentation on Replica[361] and LERF-Mask[102]. Following [100], we report mIoU and F-score on Replica, and mIoU and mBIoU on LERF-Mask.
@@
TABLE A.3: Quantitative 3D interactive segmentation on NVOS [363] and SPIn-NeRF[144] in terms of mIoU and mAcc.
@@
TABLE A.4: Performance comparison of open-vocabulary 3D semantic segmentation on ScanNet [356] benchmark. The evaluation metrics are mIoU and mAcc, respectively.
@@
$\bullet$ Metrics. Segmentation performance in 3DGS is typically evaluated from three aspects: region accuracy, boundary quality, and instance-level quality.
Region accuracy. Mean Intersection over Union (mIoU) measures the average overlap ratio between predicted and ground-truth regions across classes, and is the most commonly used metric for semantic segmentation. Mean Accuracy (mAcc) measures the proportion of correctly classified pixels or instances, reflecting overall classification correctness, but it is generally less sensitive than mIoU to class imbalance or boundary errors. F-score, typically computed under an IoU threshold (e.g., 0.5), balances precision and recall and is often used as a complementary measure for detection- or instance-style evaluation.
Boundary quality. Mean Boundary Intersection over Union (mBIoU) computes IoU over narrow boundary regions and is therefore more sensitive to contour quality than standard mIoU. This is particularly useful when evaluating fine-grained masks or object boundaries.
Instance-level quality. Panoptic Quality (PQ) [366] jointly measures recognition quality (RQ) and segmentation quality (SQ), and is more suitable when instance separation is important.
Overall, these metrics are complementary: mIoU and mAcc emphasize region correctness, mBIoU highlights boundary precision, while PQ and F-score are more informative for instance-aware settings.
$\bullet$ Results. Following the taxonomy in Sec. 2.2.1, we report results across four representative segmentation settings. For 3D instance segmentation (Table A.2), Unified-Lift [100] achieves the best overall performance on both the Replica [361] and LERF-Mask [102] datasets, demonstrating the effectiveness of consistency-aware end-to-end learning. In the setting of 3D interactive segmentation (Table A.3), GaussianCut [109] performs best on NVOS [355], while Click-Gaussian [104] attains the highest mIoU on SPIn-NeRF [144]. Notably, GaussianCut is a training-free method that flexibly supports both 2D clicks and text prompts as input. For open-vocabulary 3D semantic segmentation (Table A.4), PanoGS [197] and InstanceGaussian [110] achieve the best performance under point-based evaluation protocols, whereas CAGS [186] performs best under 3D Gaussian-based evaluation. In the open-vocabulary 2D semantic segmentation setting, VLGaussian [108] reaches a state-of-the-art mIoU of 97.1% on the 3D-OVS benchmark [140], significantly outperforming existing approaches in Table A.5. Table A.6 further presents results for two subtasks: open-vocabulary 2D localization (detection) and semantic segmentation. FMLGS [180] achieves the best performance on the segmentation task, while VLGaussian [108] leads in localization. Across these settings, most methods leverage powerful vision-language foundation models such as CLIP and SAM to enhance open-vocabulary capabilities.
TABLE A.5: Quantitative open-vocabulary 2D semantic segmentation on 3D-OVS [140] dataset in terms of mIoU.
@@
TABLE A.6: Quantitative results on LERF-OVS [106] Dataset. We report the open-vocabulary localization accuracy (mAcc) and 2D semantic segmentation (mIoU).
@@
ReLaGS[200] [CVPR’26] 51.2 - 64.7 - 81.0 - 60.6 -
TABLE A.7: Statistics of representative 3DGS editing datasets. See Sec. 4.2 for more detailed descriptions.
  Dataset Venue #Scenes #Views(Avg.) Characterization
TABLE A.8: Performance comparison of 3DGS editing on Mip-NeRF360 [357] and IN2N [152] datasets. The similarity for CLIP Text-Image Direction,CLIP Text-Image, CLIP Image-Image, DINO are denoted as $\text{CLIP}_{dir}$, $\text{CLIP}_{T2I}$, $\text{CLIP}_{I2I}$, and $\text{DINO}_{sim}$, respectively.
@@
TABLE A.9: Performance comparison of 3DGS style transfer on Mip-NeRF360 [357] dataset, evaluated by LPIPS and RMSE.
@@
A.1.2 Performance Benchmarking: 3DGS Editing
$\bullet$ Datasets. We summarize the key features of the datasets in Table A.7, with detailed descriptions provided below.
@@ A.1.2 Performance Benchmarking: 3DGS Editing
$\bullet$ Metrics. Evaluation in 3DGS editing is inherently multi-dimensional, since a successful edit should ideally preserve scene identity and geometry while also satisfying the editing instruction. Accordingly, existing metrics can be grouped into four categories: text-alignment metrics, content-preservation metrics, image-fidelity metrics, and aesthetics and human evaluation.
Text alignment. CLIP Text-Image Similarity [124] measures how well the edited image matches the target text prompt. CLIP Text-Image Direction Similarity [124] further evaluates whether the change from the original image to the edited image follows the semantic direction specified by the text instruction, making it particularly useful for instruction-guided editing.
Content preservation. CLIP Image-Image Similarity [124] and DINO Similarity [378] assess how well the edited result preserves the original subject or scene content. These metrics are helpful for measuring identity preservation, although they may not fully reflect whether the edit is semantically successful.
Image-level fidelity. Peak Signal-to-Noise Ratio (PSNR) [379], Structural Similarity (SSIM) [380], and RMSE mainly measure pixel-level fidelity with respect to reference images, while Learned Perceptual Image Patch Similarity (LPIPS) [381] is more correlated with perceptual similarity in deep feature space. Frechet Inception Distance (FID) [382] compares the distribution of generated images against real images and is more suitable for measuring overall realism at the dataset level rather than edit faithfulness for individual samples.
Aesthetics and human evaluation. Aesthetic score [383] reflects subjective visual appeal, while user study is often necessary to evaluate aspects such as instruction faithfulness, scene consistency, and user preference that may not be well captured by automatic metrics.
Overall, no single metric is sufficient for 3DGS editing. CLIP-based scores emphasize semantic alignment, image-similarity metrics favor preservation, and pixel-/perceptual metrics reflect fidelity, so most works require multiple complementary metrics for a meaningful evaluation.
@@
$\bullet$ Results. For 3D image editing tasks, the most commonly used datasets are Mip-NeRF360 [357] and IN2N [152]. We adopt these two datasets as benchmarks to record the performance of various methods. However, as most existing 3D editing methods are evaluated on different scene subsets, it remains difficult to perform a comprehensive and fair comparison under a unified protocol. Given that IN2N [152] is widely adopted as a baseline, we report the relative gains of each method over IN2N to enable an indirect yet relatively fair comparison, as summarized in Table A.8. Overall, GaussianVTON[251] demonstrates the best performance across multiple metrics. For the style transfer task, we use Mip-NeRF 360 [357] as the benchmark and compare the results of different methods in Table A.9, where SemanticSplatSty.[264] achieves the best overall performance.
TABLE A.10: Statistics of representative 3DGS generation datasets. See Sec. 4.3 for more detailed descriptions.
  Dataset Venue #Type #Scenes Characterization
[…]
@@
TABLE A.11: Quantitative 3DGS generation on GSO [359] dataset. CLIP Image-Image Similarity as $\text{CLIP}_{I2I}$.
@@
TABLE A.12: Quantitative 3DGS generation on Objaverse [358] dataset. CLIP Text-Image Similarity as $\text{CLIP}_{T2I}$.
@@
A.1.3 Performance Benchmarking: 3DGS Generation
$\bullet$ Datasets. We summarize the key features of the datasets in Table A.10, with detailed descriptions provided below.